A physics instructor performs a demonstration where a current-carrying wire is made to levitate. The wire has a mass per unit length of 0.590 g/cm and carries a current of 2.06 A horizontally to the south. Find the magnitude and direction of the minimum magnetic field needed to lift this wire vertically upward. (Enter the magnitude in T.)

Magnitude:?

Direction:?
Answer #1

as we know

mg = Bi L

(m/L) g = B i

0.59* 10^-3 / 10^-2* 9.8 = B * 2.06

B = 0.2807 T

========

field directed to the right

======

Comment before rate in case any doubt, will reply for sure.. goodluck